# Temporally Consistent Object-Centric Learning by Contrasting Slots

### v1

@@ Abstract
Unsupervised object-centric learning from videos is a promising approach to extract structured representations from large, unlabeled collections of videos. To support downstream tasks like autonomous control, these representations must be both compositional and temporally consistent. Existing approaches based on recurrent processing often lack long-term stability across frames because their training objective does not enforce temporal consistency. In this work, we introduce a novel object-level temporal contrastive loss for video object-centric models that explicitly promotes temporal consistency. Our method significantly improves the temporal consistency of the learned object-centric representations, yielding more reliable video decompositions that facilitate challenging downstream tasks such as unsupervised object dynamics prediction. Furthermore, the inductive bias added by our loss strongly improves object discovery, leading to state-of-the-art results on both synthetic and real-world datasets, outperforming even weakly-supervised methods that leverage motion masks as additional cues. Visit slotcontrast for videos and further details.
@@ 1 Introduction
Object-centric learning (OCL) [4, 13, 30, 41] is a rapidly advancing area of visual representation learning that enables autonomous systems to represent, understand, and model high-dimensional data directly in terms of its constituent entities. Structured object-centric representations (often referred to as slots [30]) facilitate generalization and robustness [8, 7] of scene representations across diverse downstream tasks, from visual question answering [1, 32, 55] to control [60, 9, 58, 17]. Of particular interest are video-based object-centric methods [27, 19, 10, 42, 2, 54, 61] that learn to represent objects that evolve and interact over time. These representations make the methods a powerful tool for applications such as unsupervised online object tracking [50, 33] and structured world modeling [26, 49, 53]. Unsupervised object-centric learning on videos has seen significant progress in recent years [2, 54, 61], mainly due to the use of pre-trained representations from self-supervised foundational models [5, 36] coupled with diverse training datasets like YouTube-VIS [56, 57]. Nevertheless, these methods still face significant challenges, especially maintaining consistent object-centric representations across time and uniquely representing each object—critical factors for successful multi-object tracking and modeling of dynamic scenes [27, 49, 53].
@@ 2 Related Work
There exists an extensive body of research [28, 19, 48, 49, 13, 22, 27, 10, 42, 46, 39, 2, 61] on discovering objects from video without any human annotations, primarily through tracking either object bounding boxes or masks. To achieve this, most of these works combine an auto-encoder framework with a simple reconstruction objective, adding inductive biases for object discovery through structured encoders [4, 30] and decoders [52]. In particular, many modern object-centric image models [41, 54, 20, 23] use a latent slot attention module [30] to extract object representations and corresponding object masks. For video data, most current methods [27, 63, 10, 42, 46, 61] connect slots across frames, with slots from the previous frame initializing those in the current frame. Notably, recent approaches [61, 38, 2] have successfully scaled object discovery to real-world unconstrained videos. To achieve this, SOLV [2] introduces temporal consistency via agglomerative clustering and prediction of middle-frame features, whereas VideoSAUR [61] learns object-centric representations by predicting temporal similarity of self-supervised features [5, 36]. While such methods can decompose short videos, they still struggle with long-term temporal consistency. In contrast, we show that learning representations that are both informative and contrastive can significantly enhance both object discovery and temporal consistency on longer videos.
@@
Achieving temporal consistency is essential for any computer vision task involving video data, whether it is tracking points, bounding boxes, segmentation masks, optical flow, or representations [31, 44, 49, 51, 47]. In object-centric learning for videos, a range of different approaches have been proposed. For example, Yu and Xu [59] apply an object-wise sequential VAE to achieve consistency; Zhao et al. [62] and Li et al. [29] use an explicit memory buffer to maintain historical slot information and a transformer as a predictor using the memory buffer to predict the future; Qian et al. [38] achieve temporal consistency by employing student-teacher distillation to establish semantic and instance correspondence over time; and Traub et al. [45] use a recurrent network with a constancy prior [16].
@@ 3.1 Semantic Recurrent Slot Attention Module
to adapt the frozen dense features for object-centric grouping. Based on the transformed encoder features $h_{t}$ and a set of slot representations of the previous timestep $S_{t-1}^{p}$, with $K$ slots $s_{t-1}^{k,p}\in S_{t-1}^{p}$, we use a recurrent grouping module to extract slot representations. The Recurrent Slot Attention module comprises a grouping module $\text{C}_{\theta}$ and a predictor module $\text{P}_{\omega}$. The former updates slot representations using the standard Slot Attention module [30] on visual features $h_{t}$ from the encoder, while the latter captures temporal and spatial interactions between slots:
@@ 3.2 Temporal Consistency through Slot Contrast
Contrastive learning is flexible in supporting diverse data sources and loss function designs. By carefully defining positive and negative examples, we can craft robust loss objectives that effectively guide self-supervised representation learning [6]. For instance, video contrastive methods like CVRL [37] leverage augmented video chunks to define positive (from the same video) and negative (from different videos) examples. We propose a novel application of a contrastive loss for temporal consistency in object-centric slot representations. In particular, we define positive samples as the representations of the same slot from two consecutive time steps within a video, while negative samples comprise all other slots across the batch between these time steps. An overview of the proposed loss is presented in Fig. 1
@@
To force each slot to be consistent in time, we aim to learn slots that are similar in time while being maximally dissimilar to other slots. Given the sets of slot representations $S_{t}$ and $S_{t+1}$ at time steps $t$ and $t+1$, we want elements $s^{i}_{t}\in S_{t}$ to be close to the next-frame slots $s^{i}_{t+1}$ corresponding to the same object, while having maximal distance to the next-frame slots $s^{k}_{t+1},k\neq i$ corresponding to other objects in the video. The corresponding InfoNCE contrastive loss [35] is defined as $\mathcal{L}_{\text{intra}}=\frac{1}{K}\sum_{i=1}^{K}\ell^{\text{\,intra}}_{i}$ with
$\ell^{\text{\,intra}}_{i}=-\log\frac{\exp(\text{sim}(s^{i}_{t},s^{i}_{t+1})/\tau)}{\sum_{k=1}^{K}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}(s^{i}_{t},s^{k}_{t+1})/\tau)},$ (4)
where $K=|S_{t+1}|$ is a number of slots per frame, $\text{sim}(u,v)=\frac{u^{\top}v}{\|u\|_{2}\|v\|_{2}}$ is the cosine similarity, $\mathbbm{1}_{[.]}$ is an indicator excluding the self-similarity of the slot $s_{i}$ from the denominator, and $\tau>0$ is a temperature parameter.
@@
$\ell_{i,j}^{\text{\,ssc}}=-\log\frac{\exp(\text{sim}(s^{i,j}_{t},s^{i,j}_{t+1})/\tau)}{\sum\limits_{b=1}^{B}\sum\limits_{k=1}^{K}\mathbbm{1}_{[k,b\neq i,j]}\exp(\text{sim}(s^{i,j}_{t},s^{k,b}_{t+1})/\tau)},$ (5)
where $B$ is a number of videos in the batch. For more details on slot-slot contrastive loss implementation, see App. C.
@@
To encourage scene decomposition we use a feature reconstruction loss, similar to DINOSAUR [41] and VideoSAUR [61]. Our final loss function combines the reconstruction loss with our proposed contrastive loss $\mathcal{L}_{\text{ssc}}$, weighted by the hyperparameter $\alpha$:
$\mathcal{L}=\sum_{t=1}^{T-1}\mathcal{L}_{\text{rec}}(\mathrm{h_{t}},\mathrm{\hat{h}_{t}})+\alpha\mathcal{L}_{\text{ssc}}(S_{t},S_{t+1}).$ (6)
@@ 4 Experiments
To evaluate our method in the controlled setting, we use MOVi-C and MOVi-E synthetic datasets generated by Kubric [15]. MOVi-C includes richly textured everyday objects, featuring up to 11 objects per scene, while MOVi-E expands this to 23 objects and introduces basic linear camera motion. In addition, to study the scalability of our method to real-world data, we evaluate our method on the real-world YouTube-VIS 2021 (YTVIS21) video dataset [57]. YTVIS21 is an unconstrained, real-world dataset sourced from YouTube, capturing a diverse range of scenes (for more details, see App. D).
@@
Similar to other object-centric video methods [10, 27, 42, 61], to evaluate consistent object discovery, we use the video foreground adjusted rand index (FG-ARI) [13], measuring how well objects are split. In addition, we evaluate the sharpness of masks using the video intersection over union with mean best overlap matching (mBO) metric [41, 61]. Both metrics are computed over the full video and thus reflect how consistent object discovery is. In addition, to investigate the effects of the temporal consistency inductive bias on the per-frame object discovery itself, we use per-frame FG-ARI (image FG-ARI), which we independently compute for each frame and average afterwards. More details can be found in App. E.
@@ 4.1 Object Discovery
We employ the DINOv2 model as our feature encoder, using ViT-S/14 for the MOVi-C dataset and ViT-B/14 for MOVi-E and YTVIS21. The slot dimension is set to 128 for MOVi-E and 64 for both MOVi-C and YTVIS21. For the MOVi datasets, we use a resolution of $(336,336)$, generating $24\!\times\!24$ patches yielding $576$ ViT tokens, while for YTVIS21, a resolution of $(518,518)$ yields $1369$ tokens. As our loss does not require full videos during training, we train our model using a sequence length of $4$ and evaluate on the full sequence length. Full implementation details are provided in App. A.
@@
We compare Slot Contrast against the previously proposed SAVi [27] and STEVE [42] that employ an image reconstruction objective and with the state-of-the-art method VideoSAUR [61] that uses self-supervised feature reconstructions. Additionally, for a fair comparison, we trained a modification of VideoSAUR with DINOv2 features (referred to as VideoSAURv2) using the same video resolution as our method. Interestingly, the default VideoSAURv2 configuration was not performing well on the simplest MOVi-C dataset (while other dataset were working better than VideoSAUR without additional changes), thus we additionally explored possible parameters options in App. F.
@@
Slot Contrast significantly outperforms both SAVi and STEVE by a wide margin. When compared to VideoSAUR using its default parameters, our approach demonstrates higher consistency in terms of video FG-ARI scores. Compared to VideoSAUR and VideoSAURv2, Slot Contrast achieves superior video scene decomposition (measured by FG-ARI scores). However, on synthetic datasets Slot Contrast’s masks are less sharp (as reflected by mBO). Notably, on the most challenging real-world YouTube-VIS data, our method surpasses both versions of VideoSAUR, achieving better performance on FG-ARI ($+6.8$) and mBO ($+4$). This shows that, given a large enough resolution and natural data inputs well aligned with DINOv2, Slot Contrast can decompose unconstrained videos into consistent object representations. More examples are illustrated in App. L.
@@
The results on MOVi-E dataset are presented in Table 2, with comparisons across additional datasets provided in App. G. Using temporal signals from the video using feature reconstruction is better than object discovery based on images. Next, additional objectives that exploit the temporal structure of the videos allow even better scene decomposition. Notably, our method, which combines a feature reconstruction objective with a simple contrastive objective, leads to state-of-the-art performance reaching $84.8$ per-frame FG-ARI, outperforming methods [3, 24] that use motion segmentation masks for object discovery.
@@
We find that the feature reconstruction baseline achieves only $16\%$ mBO vs. our method obtains $21\%$ mBO on fully occluded objects. Our results suggest that Slot Contrast significantly enhances consistency during object disappearances and reappearances. We refer the reader to Fig. 4 and App. J for visual examples and more details.
Figure 4: Comparison of the Feature Reconstruction (Feat. Rec.) baseline, the slot-slot contrastive loss using only slots from the same video as the contrast set (Intra-video Contrast), and Slot Contrast on the MOVi-C dataset. Failures and successes to track one object are marked with pink and green, respectively.
@@ 4.2 Object Dynamics Prediction
To evaluate performance on the task of predicting object dynamics, we train a dynamics module using the object-centric representations inferred by a pretrained object-centric model. For this dynamics module, we select SlotFormer [53], which predicts the slots autoregressively for $K$ rollout steps based on the slots inferred from $T$ burn-in frames preceding the prediction horizon. Since SlotFormer is trained independently from the object-centric model, we first train the latter, subsequently extending the datasets with the inferred slots for each frame. This approach avoids the computational complexity of training SlotFormer by removing the necessity to encode frames into the slot space at each training step. A brief introduction to SlotFormer and the implementation details can be found in App. I.
@@ 4.3 Analysis
In video object-centric learning, slots are typically initialized based on those from the previous time step [27], while the first frame is initialized from learnable parameters. While previous real-world object-centric methods mostly used random initialization samples from Gaussian distribution [61, 2], in this work, we study the impact of the type of initialization under our contrastive objective. The findings are outlined in Table 5. Our experiments indicate that when combined with slot-slot contrastive loss, learned initialization significantly outperforms random initialization. We hypothesize that this improvement stems from the ability of learned initializations to shape the initial state in a way that enhances contrastiveness, a benefit not achievable with random initialization. For similarity visualizations and more details, see  App. B.
@@
In models based on the Slot Attention mechanism, all available slots are typically utilized [41], leading to a mismatch between the predicted number of components in scene decomposition and the ground-truth number of objects in the scene. This can cause the random splitting of the objects between slots and non-consistent scene representations when slots are reassigned from one object to a part of another object. To address this challenge, it is important to study whether redundant slots can be effectively deactivated, especially in complex real-world videos. Recently, AdaSlot [11] introduced a discrete slot sampling module, coupled with a complexity-aware prior, to penalize redundant slots explicitly. Similarly, SOLV [2] used agglomerative clustering to merge redundant slots. In this work, we investigate whether Slot Contrast is capable of accurately determining the number of objects in a scene without relying on explicit priors to minimize the number of active slots.

### v2

@@ Abstract
Unsupervised object-centric learning from videos is a promising approach to extract structured representations from large, unlabeled collections of videos. To support downstream tasks like autonomous control, these representations must be both compositional and temporally consistent. Existing approaches based on recurrent processing often lack long-term stability across frames because their training objective does not enforce temporal consistency. In this work, we introduce a novel object-level temporal contrastive loss for video object-centric models that explicitly promotes temporal consistency. Our method significantly improves the temporal consistency of the learned object-centric representations, yielding more reliable video decompositions that facilitate challenging downstream tasks such as unsupervised object dynamics prediction. Furthermore, the inductive bias added by our loss strongly improves object discovery, leading to state-of-the-art results on both synthetic and real-world datasets, outperforming even weakly-supervised methods that leverage motion masks as additional cues.
Visit slotcontrast.github.io for videos and further details.
@@ 1 Introduction
Object-centric learning (OCL) [4, 14, 32, 44] is a rapidly advancing area of visual representation learning that enables autonomous systems to represent, understand, and model high-dimensional data directly in terms of its constituent entities. Structured object-centric representations (often referred to as slots [32]) facilitate generalization and robustness [9, 7] of scene representations across diverse downstream tasks, from visual question answering [1, 34, 58, 8] to control [63, 10, 61, 18]. Of particular interest are video-based object-centric methods [28, 20, 11, 45, 2, 57, 64] that learn to represent objects that evolve and interact over time. These representations make the methods powerful tools for applications such as unsupervised online object tracking [53, 35] and structured world modeling [27, 52, 56]. Unsupervised object-centric learning on videos has seen significant progress in recent years [2, 57, 64], mainly due to the use of pre-trained representations from self-supervised foundational models [5, 38] coupled with diverse training datasets like YouTube-VIS [59, 60]. Nevertheless, these methods still face significant challenges, especially maintaining consistent object-centric representations across time and uniquely representing each object—critical factors for successful multi-object tracking and modeling of dynamic scenes [28, 52, 56].
@@ 2 Related Work
There exists an extensive body of research [30, 20, 51, 52, 14, 23, 28, 11, 45, 49, 42, 2, 64] on discovering objects from video without any human annotations, primarily through tracking either object bounding boxes or masks. To achieve this, most of these works combine an auto-encoder framework with a simple reconstruction objective, adding inductive biases for object discovery through structured encoders [4, 32] and decoders [55]. In particular, many modern object-centric image models [44, 57, 21, 24, 8] use a latent slot attention module [32] to extract object representations and corresponding object masks. For video data, most current methods [28, 66, 11, 45, 49, 64] connect slots across frames, with slots from the previous frame initializing those in the current frame. Notably, recent approaches [64, 40, 2] have successfully scaled object discovery to real-world unconstrained videos. To achieve this, SOLV [2] introduces temporal consistency via agglomerative clustering and prediction of middle-frame features, whereas VideoSAUR [64] learns object-centric representations by predicting temporal similarity of self-supervised features [5, 38]. While such methods can decompose short videos, they still struggle with long-term temporal consistency. In contrast, we show that learning representations that are both informative and contrastive can significantly enhance both object discovery and temporal consistency on longer videos.
@@
Achieving temporal consistency is essential for any computer vision task involving video data, whether it is tracking points, bounding boxes, segmentation masks, optical flow, or representations [33, 47, 52, 54, 50, 29, 41]. In object-centric learning for videos, a range of different approaches have been proposed. For example, Yu and Xu [62] apply an object-wise sequential VAE to achieve consistency; Zhao et al. [65] and Li et al. [31] use an explicit memory buffer to maintain historical slot information and a transformer as a predictor using the memory buffer to predict the future; Qian et al. [40] achieve temporal consistency by employing student-teacher distillation to establish semantic and instance correspondence over time; and Traub et al. [48] use a recurrent network with a constancy prior [17].
@@ 3.1 Semantic Recurrent Slot Attention Module
to adapt the frozen dense features for object-centric grouping (see  App. I for more details and visualizations). Based on the transformed encoder features $h_{t}$ and a set of slot representations of the previous timestep $S_{t-1}^{p}$, with $K$ slots $s_{t-1}^{k,p}\in S_{t-1}^{p}$, we use a recurrent grouping module to extract slot representations. The Recurrent Slot Attention module comprises a grouping module $\text{C}_{\theta}$ and a predictor module $\text{P}_{\omega}$. The former updates slot representations using the standard Slot Attention module [32] on visual features $h_{t}$ from the encoder, while the latter captures temporal and spatial interactions between slots:
@@ 3.2 Temporal Consistency through Slot Contrast
Contrastive learning is flexible in supporting diverse data sources and loss function designs. By carefully defining positive and negative examples, we can craft robust loss objectives that effectively guide self-supervised representation learning [6]. For instance, video contrastive methods like CVRL [39] leverage augmented video chunks to define positive (from the same video) and negative (from different videos) examples. In object-centric learning Didolkar et al. [8] employed a contrastive loss function to gain controllability over slot representations guided by language. We propose a novel application of a contrastive loss for temporal consistency in object-centric slot representations. In particular, we define positive samples as the representations of the same slot from two consecutive time steps within a video, while negative samples comprise all other slots across the batch between these time steps. An overview of the proposed loss is presented in Fig. 2
@@
To force each slot to be consistent in time, we aim to learn slots that are similar in time while being maximally dissimilar to other slots. Given the sets of slot representations $S_{t-1}$ and $S_{t}$ at time steps $t-1$ and $t$, we want elements $s^{i}_{t-1}\in S_{t-1}$ to be close to the next-frame slots $s^{i}_{t}$ corresponding to the same object, while having maximal distance to the next-frame slots $s^{k}_{t},k\neq i$ corresponding to other objects in the video. The corresponding InfoNCE contrastive loss [37] is defined as $\mathcal{L}_{\text{intra}}=\frac{1}{K}\sum_{i=1}^{K}\ell^{\text{\,intra}}_{i}$ with
$\small\ell^{\text{\,intra}}_{i}=-\log\frac{\exp(\text{sim}(s^{i}_{t-1},s^{i}_{t})/\tau)}{\sum_{k=1}^{K}\mathbbm{1}_{[k\neq i]}\exp(\text{sim}(s^{i}_{t-1},s^{k}_{t})/\tau)},$ (4)
where $K=|S_{t}|$ is a number of slots per frame, $\text{sim}(u,v)=\frac{u^{\top}v}{\|u\|_{2}\|v\|_{2}}$ is the cosine similarity, $\mathbbm{1}_{[.]}$ is an indicator excluding the self-similarity of the slot $s_{i}$ from the denominator, and $\tau>0$ is a temperature parameter.
@@
$\small\ell_{i,j}^{\text{\,ssc}}=-\log\frac{\exp(\text{sim}(s^{i,j}_{t-1},s^{i,j}_{t})/\tau)}{\sum\limits_{b=1}^{B}\sum\limits_{k=1}^{K}\mathbbm{1}_{[k,b\neq i,j]}\exp(\text{sim}(s^{i,j}_{t-1},s^{k,b}_{t})/\tau)},$ (5)
where $B$ is a number of videos in the batch and $s^{i,j}_{t}$ denotes the $i$-th slot of the $j$-th video at time $t$. For more details on slot-slot contrastive loss implementation, see App. C.
@@
To encourage scene decomposition we use a feature reconstruction loss, similar to DINOSAUR [44] and VideoSAUR [64]. Our final loss function combines the reconstruction loss with our proposed contrastive loss $\mathcal{L}_{\text{ssc}}$, weighted by the hyperparameter $\alpha$ (see Table LABEL:tab:implementation_details for details on how the hyperparameters are set):
$\small\mathcal{L}=\sum_{t=1}^{T-1}\mathcal{L}_{\text{rec}}(\mathrm{h_{t}},\mathrm{\hat{h}_{t}})+\alpha\mathcal{L}_{\text{ssc}}(S_{t-1},S_{t}).$ (6)
@@ 4 Experiments
To evaluate our method in the controlled setting, we use MOVi-C and MOVi-E synthetic datasets generated by Kubric [16]. MOVi-C includes richly textured everyday objects, featuring up to 11 objects per scene, while MOVi-E expands this to 23 objects and introduces basic linear camera motion. In addition, to study the scalability of our method to real-world data, we evaluate our method on the real-world YouTube-VIS 2021 (YTVIS21) video dataset [60]. YTVIS21 is an unconstrained, real-world dataset sourced from YouTube, capturing a diverse range of scenes (for more details, see App. E).
@@
Similar to other object-centric video methods [11, 28, 45, 64], to evaluate consistent object discovery, we use the video foreground adjusted rand index (FG-ARI) [14], measuring how well objects are split. In addition, we evaluate the sharpness of masks using the video intersection over union with mean best overlap matching (mBO) metric [44, 64]. Both metrics are computed over the full video and thus reflect how consistent object discovery is. In addition, to investigate the effects of the temporal consistency inductive bias on the per-frame object discovery itself, we use per-frame FG-ARI (image FG-ARI), which we independently compute for each frame and average afterwards. More details can be found in App. F.
@@ 4.1 Object Discovery
We employ the DINOv2 model as our feature encoder, using ViT-S/14 for the MOVi-C dataset and ViT-B/14 for MOVi-E and YTVIS21. The slot dimension is set to 128 for MOVi-E and 64 for both MOVi-C and YTVIS21. For the MOVi datasets, we use a resolution of $(336,336)$, generating $24\!\times\!24$ patches yielding $576$ ViT tokens, while for YTVIS21, a resolution of $(518,518)$ yields $1369$ tokens. Full details are provided in App. A.
@@
We compare Slot Contrast against the previously proposed SAVi [28] and STEVE [45] that employ an image reconstruction objective and with the state-of-the-art method VideoSAUR [64] that uses self-supervised feature reconstructions. Additionally, for a fair comparison, we trained a modification of VideoSAUR with DINOv2 features (referred to as VideoSAURv2). The implementation details are provided in App. G. In addition, to assess how closely Slot Contrast approaches supervised methods, we compared it with SAM2 [41] as a supervised zero-shot baseline for temporal consistency and to weakly-supervised by depth SAVi++ [11] method. The results are in App. G.
@@
Slot Contrast significantly outperforms both SAVi and STEVE by a wide margin. When compared to VideoSAUR using its default parameters, our approach demonstrates higher consistency in terms of video FG-ARI scores. Compared to VideoSAUR and VideoSAURv2, Slot Contrast achieves superior video scene decomposition (measured by FG-ARI scores). However, on synthetic datasets Slot Contrast’s masks are less sharp (as reflected by mBO). Notably, on the most challenging real-world YouTube-VIS data, our method surpasses both versions of VideoSAUR, achieving better performance on FG-ARI ($+6.8$) and mBO ($+4$). This shows that, given a large enough resolution and natural data inputs well aligned with DINOv2, Slot Contrast can decompose unconstrained videos into consistent object representations. More examples are illustrated in App. M.
@@
The results on MOVi-E dataset are presented in Table 2, with comparisons across additional datasets provided in App. H. Using temporal signals from the video using feature reconstruction is better than object discovery based on images. Next, additional objectives that exploit the temporal structure of the videos allow even better scene decomposition. Notably, our method, which combines a feature reconstruction objective with a simple contrastive objective, leads to state-of-the-art performance reaching $84.8$ per-frame FG-ARI, outperforming methods [3, 25] that use motion segmentation masks for object discovery.
@@
We find that the feature reconstruction baseline achieves only $16\%$ mBO vs. our method obtains $21\%$ mBO on fully occluded objects. Our results suggest that Slot Contrast significantly enhances consistency during object disappearances and reappearances. We refer the reader to Fig. 4 and App. K for visual examples and more details.
Figure 4: Comparison of the Feature Reconstruction (Feat. Rec.) baseline, the slot-slot contrastive loss using only slots from the same video as the contrast set (Intra-video Contrast), and Slot Contrast on the MOVi-C dataset.
@@ 4.2 Object Dynamics Prediction
To evaluate performance on the task of predicting object dynamics, we train a dynamics module using the object-centric representations inferred by a pretrained object-centric model. For this dynamics module, we select SlotFormer [56], which predicts the slots autoregressively for $K$ rollout steps based on the slots inferred from $T$ burn-in frames preceding the prediction horizon. In our setup, we use 14 burn-in frames and 10 rollout steps. Since SlotFormer is trained independently from the object-centric model, we first train the latter, subsequently extending the datasets with the inferred slots for each frame. This approach avoids the computational complexity of training SlotFormer by removing the necessity to encode frames into the slot space at each training step. A brief introduction to SlotFormer and the implementation details can be found in App. J.
@@ 4.3 Analysis
In video object-centric learning, slots are typically initialized based on those from the previous time step [28], while the first frame is initialized from learnable parameters. Previous real-world object-centric methods mostly used random initialization samples from Gaussian distribution [64, 2], in this work, we study the impact of the type of initialization under our contrastive objective. The findings are outlined in Table 5. Our experiments indicate that when combined with slot-slot contrastive loss, learned initialization significantly outperforms random initialization. We hypothesize that this improvement stems from the ability of learned initializations to shape the initial state in a way that enhances contrastiveness, a benefit not achievable with random initialization. For similarity visualizations and more details, see  App. B.
@@
In models based on the Slot Attention mechanism, all available slots are typically utilized [44], leading to a mismatch between the predicted number of components in scene decomposition and the ground-truth number of objects in the scene. This can cause the random splitting of the objects between slots and non-consistent scene representations when slots are reassigned from one object to a part of another object. To address this challenge, it is important to study whether redundant slots can be effectively deactivated. Recently, AdaSlot [12] introduced a discrete slot sampling module, coupled with a complexity-aware prior, to penalize redundant slots explicitly. Similarly, SOLV [2] used agglomerative clustering to merge redundant slots. In this work, we investigate whether Slot Contrast is capable of accurately determining the number of objects in a scene without relying on explicit priors to minimize the number of active slots.